If the following solutions are mixed, is the resulting solution acidic, basic, or neutral?

1)65.0 mL of 0.0500 M HClO4 and 40.0 mL of 0.0750 M NaOH

2)125.0 mL of 0.100 M HNO3 and 80.0 mL of 0.0750 M Ca (OH) 2
Answer #1

1)

Mol of HClO4 added = M(HClO4)*V(HClO4)

= 0.0500 M * 65.0 mL

= 3.25 mmol

Mol of NaOH added = M(NaOH)*V(NaOH)

= 0.0750 M * 40.0 mL

= 3.00 mmol

Mol of H+ is more than mol of OH-.

So, the solution would be acidic

Answer: Acidic

2)

Mol of HNO3 added = M(HNO3)*V(HNO3)

= 0.100 M * 125.0 mL

= 12.5 mmol

Mol of Ca(OH)2 added = M(Ca(OH)2)*V(Ca(OH)2)

= 0.0750 M * 80.0 mL

= 6.0 mmol

1 mol of Ca(OH)2 has 2 mol of OH-

So,

Mol of OH- = 2*6.0 mmol = 12 mmol

Mol of H+ is more than mol of OH-.

So, the solution would be acidic

Answer: Acidic